# Tunable topological phase in 2D ScV6Sn6 kagome material

### v1

@@ I INTRODUCTION
The study of exotic phenomena in kagome materials has progressed significantly with the development of topological quantum matter theories [1]. These advances have facilitated the discovery of various novel topological materials, including topological insulators (TIs) and topological semimetals (TSMs). Both classes exhibit unique features, such as insulating bulk states and boundary states protected by symmetries [2, 3, 4, 5]. These boundary states often consist of Dirac or Weyl points, which connect energy bands of opposite parity. In the presence of spin-orbit coupling (SOC), the degeneracy at these points can be lifted, forming a topological gap. Notably, these intrinsic properties do not require external fields to manifest.
@@
Nonmagnetic TIs and TSMs are characterized by topological states protected by time-reversal symmetry (TRS), while magnetic TSMs exhibit unique phases, such as Weyl nodes and Chern numbers, arising from the breaking of crystal symmetries [6]. Despite significant advances in understanding nonmagnetic topological phases, both insulating and (semi)metallic—magnetic topological materials remain relatively less explored. The complex interactions within these materials pose a greater challenge for theoretical predictions compared to their nonmagnetic counterparts. However, they hold considerable technological potential due to the additional degrees of freedom provided by magnetism, which enable novel ways of manipulating topological states. Recent theoretical and experimental advances have deepened our understanding of magnetic topological materials [7, 8, 9].
@@
A key topological invariant in such systems is the Chern number (C), which breaks TRS while preserving particle-hole symmetry. The Chern number is given by the integral of the Berry curvature over the Brillouin zone and corresponds to the number of stable chiral modes in the occupied bulk states [10]. In two-dimensional (2D) materials, a non-trivial value of C indicates a 2D Chern insulator, characterized by delocalized boundary states at the edges of the material [11, 12]. Another important topological invariant is the $\mathbb{Z}_{2}$ invariant, denoted as $(v_{0},v_{1},v_{2},v_{3})$, used to classify both 2D and three-dimensional (3D) topological insulators [1]. The $\mathbb{Z}_{2}$ invariant distinguishes trivial insulators ($\mathbb{Z}_{2}=0$) from topological insulator states ($\mathbb{Z}_{2}=1$), with a nontrivial $\mathbb{Z}_{2}$ associated with the quantum spin Hall (QSH) phase. The topological phase is classified as weak ($v_{0}=0$) or strong ($v_{0}=1$), depending on the value of $v_{0}$ [12].
@@
Recent studies have predicted strong nonmagnetic topological insulator phases in kagome materials, which are known for their complex electronic properties, including symmetry-breaking charge density waves (CDWs) and intricate phase diagrams [13, 14, 15, 16]. For instance, materials such as GdV6Sn6 and YV6Sn6 demonstrate behavior akin to topological insulators in their paramagnetic states [17, 16]. Similarly, bulk TbV6Ge6, a member of the RV6Ge6 family, has been predicted to exhibit topological crystalline insulator properties, with these topological features protected by its structural symmetry [12]. In the realm of 2D materials, the study of magnetic topological insulators and semimetals is just beginning, with only a few examples, such as MnTe(Bi2Te3) and MnBi2Te4, identified to date. MnBi2Te4 exhibits antiferromagnetic behavior, with a Néel temperature (TN) of approximately 25 K [9]. Other notable discoveries include the kagome ferromagnetic (FM) Weyl semimetal Co3Sn2S2 and the magnetically induced Weyl semimetal GdPtBi [18, 19].
@@
A series of magnetic topological semimetals with significant experimental potential have been predicted and identified, driven by searches for large anomalous Hall effects (AHE) and ab initio calculations [20, 21]. The anomalous Hall effect, first discovered by Edwin Hall, occurs in all ferromagnetic semimetals and metals, where the Hall resistivity’s response to an applied magnetic field mirrors the magnetization response [22]. The AHE is now understood to be significantly influenced by the Berry curvature, which acts like a magnetic monopole at energy level crossings, also known as Weyl points [23]. While most studies on TIs and TSMs have focused on intrinsic topological properties, recent research has explored the tunability of these phases through external perturbations. For example, applying external electric fields (E-fields) can shift the position of Dirac cones in momentum space, inducing topological phase transitions. In bilayer Bi(111), an infinitesimal E-field can transition the material from a non-trivial to a trivial state [24]. Similarly, phosphorene, a 2D material derived from black phosphorus, can transition from a normal insulator to a topological insulator and eventually to a metal under an applied electric field [25]. These findings underscore the potential for controlling topological phases through external fields.
@@
Motivated by the emerging topological properties that characterize low-dimensional systems, we investigated the electronic structure and topological features of the two-dimensional kagome material ScV6Sn6, a member of the HfFe6Ge6 family. In its bulk form, ScV6Sn6 is a nontrivial metal exhibiting a tunable charge density wave (CDW) state with a first-order phase transition and a propagation vector of $\left(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3}\right)$, as determined from x-ray and neutron scattering experiments [26, 27]. In this work, we explore the emergent properties of its 2D counterpart. Our ab initio simulations indicate that 2D ScV6Sn6 is a topological material with a ferromagnetic ground state characterized by a mean-field Curie temperature of $T_{C}\approx 89\,$K and a magnetic moment of approximately $0.86\,\mu_{B}$ per atom. This behavior contrasts with the nonmagnetic nature of its three-dimensional bulk form. The material is further characterized by Weyl points formed by pairs of topological bands near the Fermi energy, with a Chern number of $|C|=1$. Furthermore, a significant anomalous Hall conductivity is observed at the Fermi level, related to the ferromagnetic ordering in the system [28, 29, 30]. These chiral states, which break time-reversal symmetry, classify 2D ScV6Sn6 as a magnetic Weyl semimetal with nontrivial topology. Additionally, we demonstrate that the nontrivial topological phase remains robust under the application of moderate electric fields, but transitions from a nontrivial to a trivial phase at an electric field strength of $\approx 0.4\,$eV/Å.
@@
Figure 1: (a) Side and (b) top views of the crystal structure of 2D ScV6Sn6. (c) Electronic band structure with the spin-up states (black bubbles) and spin-down states (red triangles). (d) Projected density of states onto atomic orbitals. The blue, red, and green lines represent the contributions from Sn, V, and Sc atoms, respectively, while the gray shading denotes the total density of states for the 2D structure. The left panel of (d) shows spin-up states, and the right panel shows spin-down states. The horizontal dotted line indicates the Fermi energy.
@@ II COMPUTATIONAL METHODS
The bulk structure of ScV6Sn6 crystallizes in the hexagonal space group $P6/mmm$ (No. 183) and comprises two layers of vanadium (V) kagome nets, each formed by six V atoms. To investigate its two-dimensional (2D) properties, a monolayer was cleaved from the bulk structure, and a vacuum layer of approximately 23 Å  was introduced along the c-axis to eliminate interactions between periodic images. The resulting 2D structure preserves the hexagonal symmetry of the bulk, with lattice constants $a=b=5.31$  (see Figure 1(a)-(b)). Density functional theory (DFT) [31] calculations were performed using the Vienna Ab initio Simulation Package (VASP) [32, 33], employing the generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [34]. Structural relaxation was carried out until the total energy and atomic forces converged to within $10^{-7}$ eV and 0.02 eV/Å, respectively. A $\Gamma$-centered $5\times 5\times 1$ k-point mesh and a plane-wave energy cutoff of 500 eV were used for these calculations. For the electronic structure analysis, we used the first-principles, spin-polarized DFT plus effective Coulomb interaction (DFT+U) including SOC to account for relativistic effects. We used a Hubbard U parameter of 6.0 eV for both V and Sc atoms, in accordance with previous studies on related systems [35, 36]. To explore the topological properties of 2D ScV6Sn6, maximally localized Wannier functions (MLWFs) [37, 11] were employed to construct a low-energy effective Hamiltonian using the Wannier90 code [38]. The Wannierization process included Sc-$d$, V-$d$, and Sn-$p$ orbitals, as they contribute significantly to the electronic states near the Fermi level. Topological characteristics, including Chern numbers and edge states, were computed using the WannierTools package [39]. A semi-infinite system was modeled with a surface slab of $\approx$250 Å  thick, and the surface Green’s function was calculated to derive the local density of states and edge state dispersion. The robustness of the topological phase was further examined by applying an external electric field. The intrinsic anomalous Hall conductivity (AHC) was computed by integrating the Berry curvature over the entire Brillouin zone for all occupied bands, offering insight into the system’s topological response under varying conditions.
@@ III RESULTS AND DISCUSSION
We investigate the electronic properties of 2D ScV6Sn6 using spin-polarized electronic band structure and density of states calculations, as shown in Figure 1 (c)-(d), obtained via DFT+U. The results confirm that 2D ScV6Sn6 is metallic, characterized by multiple Weyl-like crossings near the Fermi level. Notably, the states around the Fermi level are predominantly contributed by V-$d$ orbitals, with smaller yet significant contributions from Sn and Sc atoms. This observation aligns with previous findings for the bulk structure [12, 48]. Our spin-polarized calculations indicate that the ferromagnetic configuration is the ground state, with a Curie temperature $T_{c}\approx 89~{}\text{K}$, estimated using a mean-field approximation. This suggests that the system exhibits broken time-reversal symmetry below $T_{c}$ due to magnetic ordering, in contrast to the nonmagnetic behavior observed in bulk ScV6Sn6 [49, 50].
@@
Figure 2: (a) Band structure of the 2D ScV6Sn6666 slab showing momentum-resolved contributions. (b) Edge state dispersion along the [110] edge, with red highlights indicating larger contributions of electronic states and blue highlights indicating smaller contributions. (c) Intrinsic anomalous Hall conductivity (AHC), $\sigma{xy}$ (in $\Omega^{-1}\text{cm}^{-1}$), as a function of energy for 2D ScV6Sn6.
@@
To explore the topological states of 2D ScV6Sn6, we focus on the edge states near the Fermi energy ($E_{f}=0.42$ eV) arising from three gapped Weyl points (labeled 1-3 in Figure 2 (a)) and the SOC-induced gap at the Fermi energy. These gaps belong to bands 55, 56, and 57, with energies of 33.0, 32.0, and 24.0 meV, respectively, with respect to the Fermi energy. We note that a Weyl point is only useful if it is located near the Fermi level, as only states at or near the Fermi energy can exhibit observable effects. Otherwise, contributions from the Weyl nodes may be obscured by other states. In the case of 2D ScV6Sn6, all the three gaps are positioned near the Fermi energy. Notably, Weyl gap 1 is exactly at the Fermi energy, while gaps 2 and 3 are located at the high symmetry K and M points. Our edge state plot in Figure 2(b) shows high intensity in regions originally gapped by SOC in the bulk band structure, indicating electronic states and band crossings. Valence and conduction band crossings at these Weyl points are clearly observed in the edge state band structure (Figure 2(b)). The calculated topological invariant, C, in the absence of external fields, is finite for the three bands, confirming the material as topologically nontrivial with three chiral modes terminating at the edge. Based on this analysis, we classify the ground state of 2D ScV6Sn6 as a Weyl semimetal with edge states near the Fermi energy, due to conduction-valence band crossings at discrete momentum points. [17, 51]. Since TRS is inherently broken, the stability of the chiral edge states depends solely on the broken TRS and the lattice geometry. The weyl nodes appear in pairs of opposite chirality and are well separated in momentum space, to prevent annihilation and ensure their stability. In the presence of SOC, 2D ScV6Sn6 exhibits a quantum anomalous Hall behavior, with the interplay between SOC and magnetic ordering offering additional stability to the chiral modes. For example, Fe3Sn2, a ferromagnetic kagome lattice, exhibits nontrivial topology due to strong SOC and magnetic order [52]. Similarly, 2D ScV6Sn6 retains nontrivial topological characteristics protected by the kagome lattice structure and the strong interplay between SOC and intrinsic magnetic ordering. These features are comparable to those observed in the ferromagnetic Weyl semimetal Co3Sn2S2 [53].
In ferromagnetic materials, the total anomalous Hall effect (AHE) conductivity, $\sigma_{H}^{A}$, arises from three main contributions: the intrinsic contribution $\sigma_{int}$, skew scattering $\sigma_{sc}$, and the side-jump mechanism $\sigma_{sj}$ [54, 55]. Mathematically, $\sigma_{H}^{A}$ is expressed as $\sigma_{H}^{A}=\sigma_{int}+\sigma_{sc}+\sigma_{sj}$. The intrinsic contribution, $\sigma_{int}$, depends on the material’s band structure and is independent of scattering rates, while $\sigma_{sc}$ and $\sigma_{sj}$ originate from extrinsic scattering processes caused by impurities or disorder. We calculated the energy-dependent intrinsic AHC ($\sigma_{int}$) using the Berry curvature derived from band calculations. Figure 2(d) shows the AHC plot, with a peak near the Fermi energy, reaching 257 $\Omega^{-1}\text{cm}^{-1}$. The strong SOC and ferromagnetic ordering in 2D ScV6Sn6 enhances the Berry curvature, which acts as an effective magnetic field in momentum space, contributing to the large AHC. The peak at $\sim$0.33 eV near the Fermi energy can be attributed to the proximity of band crossings to the Fermi energy. The large anomalous Hall effect is believed to arise from the Berry curvature hotspots near the Fermi level. SOC-induced small gaps act as these Berry curvature hotspots, which in turn lead to the observed AHC. Currently, there are limited reports on AHC in 2D materials, making our findings particularly significant. This study provides valuable insights into the anomalous Hall effect in 2D kagome materials and contributes to the broader understanding of topological phenomena in these systems.
Previous studies on 2D ferromagnetic Dirac materials have reported intrinsic anomalous Hall conductivity (AHC) values reaching up to approximately 300 $\Omega^{-1}\text{cm}^{-1}$, driven by strong SOC and broken TRS [56]. In Fe3GeTe2 monolayers, an intrinsic AHC of $\approx$350 $\Omega^{-1}\text{cm}^{-1}$ has been reported, with further enhancement achieved through external tuning methods such as ionic-liquid gating [57]. Similarly, studies on MnBi2Te4 have observed intrinsic AHC values ranging from 250 to 300 $\Omega^{-1}\text{cm}^{-1}$, attributed to contributions from Berry curvature and canted antiferromagnetic ordering. To the best of our knowledge, there are no previous reports on AHC in 2D kagome materials. Our predicted intrinsic AHC value of $\approx$257 $\Omega^{-1}\text{cm}^{-1}$ falls within the range of values reported for other 2D systems. Moreover, it is consistent with the values observed in bulk kagome materials, which range from $\sim$380 $\Omega^{-1}\text{cm}^{-1}$ for bulk LiMn6Sn6 [54] to as high as 500 $\Omega^{-1}\text{cm}^{-1}$ for Fe3Sn2 [58].
rev

### v2

@@ I INTRODUCTION
The study of exotic phenomena in kagome materials has progressed significantly with the development of topological quantum matter theories hasan2015topological . These advances have facilitated the discovery of various novel topological materials, including topological insulators (TIs) and topological semimetals (TSMs). Both classes exhibit unique features, such as insulating bulk states and boundary states protected by symmetries RevModPhys.82.3045 ; kane2005quantum ; niu2015two ; wang2013organic . These boundary states often consist of Dirac or Weyl points, which connect energy bands of opposite parity. In the presence of spin-orbit coupling (SOC), the degeneracy at these points can be lifted, forming a topological gap. Notably, these intrinsic properties do not require external fields to manifest.
@@
Nonmagnetic TIs and TSMs are characterized by topological states protected by time-reversal symmetry (TRS), while magnetic TSMs exhibit unique phases, such as Weyl nodes and Chern numbers, arising from the breaking of crystal symmetries okada2013observation . Despite significant advances in understanding nonmagnetic topological phases, both insulating and (semi)metallic—magnetic topological materials remain relatively less explored. The complex interactions within these materials pose a greater challenge for theoretical predictions compared to their nonmagnetic counterparts. However, they hold considerable technological potential due to the additional degrees of freedom provided by magnetism, which enable novel ways of manipulating topological states. Recent theoretical and experimental advances have deepened our understanding of magnetic topological materials xu2020high ; watanabe2018structure ; otrokov2019prediction .
@@
A key topological invariant in such systems is the Chern number (C), which breaks TRS while preserving particle-hole symmetry. The Chern number is given by the integral of the Berry curvature over the Brillouin zone and corresponds to the number of stable chiral modes in the occupied bulk states sheng2006quantum . In two-dimensional (2D) materials, a non-trivial value of C indicates a 2D Chern insulator, characterized by delocalized boundary states at the edges of the material prodan2010entanglement ; ahsan2023prediction . Another important topological invariant is the $\mathbb{Z}_{2}$ invariant, denoted as $(v_{0},v_{1},v_{2},v_{3})$, used to classify both 2D and three-dimensional (3D) topological insulators hasan2015topological . The $\mathbb{Z}_{2}$ invariant distinguishes trivial insulators ($\mathbb{Z}_{2}=0$) from topological insulator states ($\mathbb{Z}_{2}=1$), with a nontrivial $\mathbb{Z}_{2}$ associated with the quantum spin Hall (QSH) phase. The topological phase is classified as weak ($v_{0}=0$) or strong ($v_{0}=1$), depending on the value of $v_{0}$ ahsan2023prediction .
@@
Recent studies have predicted strong nonmagnetic topological insulator phases in kagome materials, which are known for their complex electronic properties, including symmetry-breaking charge density waves (CDWs) and intricate phase diagrams teng2022discovery ; tan2021charge ; yu2021concurrence ; hu2022tunable . For instance, materials such as GdV6Sn6 and YV6Sn6 demonstrate behavior akin to topological insulators in their paramagnetic states pokharel2021electronic ; hu2022tunable . Similarly, bulk TbV6Ge6, a member of the RV6Ge6 family, has been predicted to exhibit topological crystalline insulator properties, with these topological features protected by its structural symmetry ahsan2023prediction . In the realm of 2D materials, the study of magnetic topological insulators and semimetals is just beginning, with only a few examples, such as MnTe(Bi2Te3) and MnBi2Te4, identified to date. MnBi2Te4 exhibits antiferromagnetic behavior, with a Néel temperature (TN) of approximately 25 K otrokov2019prediction . Other notable discoveries include the kagome ferromagnetic (FM) Weyl semimetal Co3Sn2S2 and the magnetically induced Weyl semimetal GdPtBi morali2019fermi ; hirschberger2016chiral .
@@
A series of magnetic topological semimetals with significant experimental potential have been predicted and identified, driven by searches for large anomalous Hall effects (AHE) and ab initio calculations kubler2014non ; zhang2017strong . The anomalous Hall effect, first discovered by Edwin Hall, occurs in all ferromagnetic semimetals and metals, where the Hall resistivity’s response to an applied magnetic field mirrors the magnetization response aliev2019novel . The AHE is now understood to be significantly influenced by the Berry curvature, which acts like a magnetic monopole at energy level crossings, also known as Weyl points Berry1984quantal . While most studies on TIs and TSMs have focused on intrinsic topological properties, recent research has explored the tunability of these phases through external perturbations. For example, applying external electric fields (E-fields) can shift the position of Dirac cones in momentum space, inducing topological phase transitions. In bilayer Bi(111), an infinitesimal E-field can transition the material from a non-trivial to a trivial state Sawahata_2019 . Similarly, phosphorene, a 2D material derived from black phosphorus, can transition from a normal insulator to a topological insulator and eventually to a metal under an applied electric field liu2015switching . Other studies have also identified tunable topological phases in 2D magnetic materials. For example, monolayer MoO which exhibits anomalous Hall states under shear strain, and single-layer RuClBr which hosts a quantum anomalous valley Hall phase with chiral spin-valley locking sun2022valley ; wu2023quantum . These findings underscore the potential of 2D materials to host controllable topological phases through external fields.
@@
Motivated by the emerging topological properties that characterize low-dimensional systems, we investigated the electronic structure and topological features of the two-dimensional kagome material ScV6Sn6, a member of the HfFe6Ge6 family. In its bulk form, ScV6Sn6 is a nontrivial metal exhibiting a tunable charge density wave (CDW) state with a first-order phase transition and a propagation vector of $\left(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3}\right)$, as determined from x-ray and neutron scattering experiments arachchige2022charge ; destefano2023pseudogap . In this work, we explore the emergent properties of its 2D counterpart. Our ab initio simulations indicate that 2D ScV6Sn6 is a topological material with a ferromagnetic ground state characterized by a mean-field Curie temperature of $T_{C}\approx 89\,$K and a magnetic moment of approximately $0.86\,\mu_{B}$ per atom. This behavior contrasts with the nonmagnetic nature of its three-dimensional bulk form. The material is further characterized by Weyl points formed by pairs of topological bands near the Fermi energy, with a Chern number of $|C|=1$. Furthermore, a significant anomalous Hall conductivity is observed at the Fermi level, related to the ferromagnetic ordering in the system chang2023colloquium ; vzelezny2023high ; ye2018massive . These chiral states, which break time-reversal symmetry, classify 2D ScV6Sn6 as a magnetic Weyl semimetal with nontrivial topology. Additionally, we demonstrate that the nontrivial topological phase remains robust under the application of moderate electric fields, but transitions from a nontrivial to a trivial phase at an electric field strength of $\approx 0.4\,$eV/Å.
@@
Figure 1: (a) Side and (b) top views of the crystal structure of 2D ScV6Sn6. (c) Electronic band structure with the spin-up states (black bubbles) and spin-down states (red triangles). (d) Projected density of states onto atomic orbitals. The blue, red, and green lines represent the contributions from Sn, V, and Sc atoms, respectively, while the gray shading denotes the total density of states for the 2D structure. The left panel of (d) shows spin-up states, and the right panel shows spin-down states. The horizontal dotted line indicates the Fermi energy at $\approx 0.40\,$ eV.
@@ II COMPUTATIONAL METHODS
The bulk structure of ScV6Sn6 crystallizes in the hexagonal space group $P6/mmm$ (No. 183) and comprises two layers of vanadium (V) kagome nets, each formed by six V atoms. To investigate its two-dimensional (2D) properties, a monolayer was cleaved from the bulk structure, and a vacuum layer of approximately 23 Å  was introduced along the c-axis to eliminate interactions between periodic images. The resulting 2D structure preserves the hexagonal symmetry of the bulk, with lattice constants $a=b=5.31$ Å  (see Figure 1(a)-(b)). Density functional theory (DFT) kohn1996density calculations were performed using the Vienna Ab initio Simulation Package (VASP) kresse1996efficiency ; kresse1996efficient , employing the generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional perdew1996generalized . Structural relaxation was carried out until the total energy and atomic forces converged to $\approx 10^{-7}$ eV and 0.02 eV/Å, respectively. A $\Gamma$-centered $5\times 5\times 1$ k-point mesh and a plane-wave energy cutoff of 500 eV were used for these calculations. For the electronic structure analysis, we employed first-principles, spin-polarized DFT calculations with and without an effective on-site Coulomb interaction (DFT+U), including spin–orbit coupling (SOC) to capture relativistic effects. To assess the role of electronic correlations, we varied the Hubbard U parameter from 0 to 6.0 eV for both V and Sc atoms, with the upper bound chosen in accordance with previous studies on related systems anisimov2005full ; BERDECIA2024113067 . To explore the topological properties of 2D ScV6Sn6, maximally localized Wannier functions (MLWFs) Marzari_2012 ; prodan2010entanglement were employed to construct a low-energy effective Hamiltonian using the Wannier90 code Pizzi2020 . The Wannierization process included Sc-$d$, V-$d$, and Sn-$p$ orbitals, as they contribute significantly to the electronic states near the Fermi level. Topological characteristics, including Chern numbers and edge states, were computed using the WannierTools package WU2017 . A semi-infinite system was modeled with a surface slab of $\approx$250 Å  thick, to ensure that the top and bottom surfaces are sufficiently decoupled, preventing artificial hybridization between edge states on opposite sides. Then the surface Green’s function was calculated to derive the local density of states and edge state dispersion. The robustness of the topological phase was further examined by applying an external electric field. The intrinsic anomalous Hall conductivity (AHC) was computed by integrating the Berry curvature over the entire Brillouin zone for all occupied bands, offering insight into the system’s topological response under varying conditions.
@@ III RESULTS AND DISCUSSION
We begin our analysis by establishing the stability of 2D ScV6Sn6, evaluating its formation energy ($E_{form}$), cleavage energy ($E_{cl}$), and mechanical and dynamical stability. The formation energy, calculated using $E_{form}=E_{tot}-\sum n\epsilon_{i}$, where $E_{tot}$ is the ground-state energy of the optimized 2D ScV6Sn6 and $\epsilon_{i}$ are the ground-state energies of the constituent elements in their bulk states, is $E_{form}\approx-0.17~\text{eV}$. This exothermic value indicates that the 2D phase is stable relative to the elemental state. The cleavage energy, calculated as $E_{cl}=(E_{l}-E_{bulk})/A$, where $E_{l}$ and $E_{bulk}$ are the energies of the cleaved layer and bulk structure, respectively, and $A$ is the surface area, was found to be $E_{cl}\approx 0.27~\text{J/m}^{2}$. This value is comparable to those of other experimentally synthesized 2D materials, such as MoS2 ($0.09-0.22~\text{J/m}^{2}$), graphene ($0.19-0.72~\text{J/m}^{2}$), and significantly lower than that of black phosphorus ($\approx 0.37~\text{J/m}^{2}$), suggesting weak interlayer interactions and high experimental synthesizability wang2015measurement ; coleman2011two ; tang2014nanomechanical ; gaur2014surface .
@@
We investigate the electronic properties of 2D ScV6Sn6 by performing spin-polarized electronic structure calculations, including band structure and density of states analyses, as shown in Figure 1(c)–(d). The results demonstrate that 2D ScV6Sn6 is metallic, exhibiting multiple Weyl-like band crossings near the Fermi energy. All topological calculations presented were performed without the inclusion of a Hubbard $U$ correction. To verify the robustness of the observed electronic and topological features, we subsequently varied the Hubbard $U$ parameter applied to the V 3d and Sc orbitals from 0 to 6 eV (see Supplementary Figure supp ). This analysis confirms that while a finite $U$ stabilizes a ferromagnetic ground state, the essential features of the electronic structure, particularly the metallic character and the Weyl-like crossings, remain qualitatively unchanged across this range. This robustness underscores the intrinsic nature of the topological metallic state in 2D ScV6Sn6.
Notably, the electronic states near the Fermi level are primarily derived from the V 3d orbitals, with secondary yet non-negligible contributions from the Sn and Sc atoms. This orbital character is consistent with prior studies on the bulk counterpart of ScV6Sn6 ahsan2023prediction ; di2023electronic . Our spin-polarized calculations further reveal that the ferromagnetic configuration is the ground state, with an estimated Curie temperature of $T_{c}\approx 89~\text{K}$, obtained via mean-field approximation. This result implies that the monolayer system undergoes spontaneous magnetic ordering and breaks time-reversal symmetry below $T_{c}$, in contrast to the nonmagnetic behavior reported for the bulk compound yi2024quantum ; yi2024tuning . To further elucidate the nature of magnetic anisotropy, we computed the magnetocrystalline anisotropy energy (MAE), defined as $\mathrm{MAE}=E_{\parallel}-E_{\perp}$, where $E_{\parallel}$ and $E_{\perp}$ are the total energies corresponding to in-plane and out-of-plane magnetization directions, respectively. The calculated MAE of 2.01 meV per formula unit is positive, indicating a lower energy for out-of-plane spin alignment. This establishes that the magnetic easy axis is oriented perpendicular to the plane, a feature that could play a critical role in stabilizing topologically nontrivial spin textures in the 2D system.
@@
Figure 2: (a) Spin–orbit-coupling (SOC)–resolved electronic band structure of the 2D ScV6Sn6 slab, showing momentum-resolved orbital contributions. (b) Edge state dispersion along the [110] edge, with red regions indicating higher spectral weight and blue indicating lower weight. (c) Intrinsic anomalous Hall conductivity (AHC), $\sigma_{xy}$ (in $\Omega^{-1}~\text{cm}^{-1}$), as a function of energy. The horizontal dashed lines in (a) and (b), and the vertical dashed line in (c), indicate the Fermi energy at $\approx 0.42\,$eV. The horizontal dashed line in (c) marks the corresponding value of $\sigma_{xy}\sim 257\,\Omega^{-1}~\text{cm}^{-1}$ at the Fermi energy.
@@
To explore the topological features of 2D ScV6Sn6, we analyze the edge states in the vicinity of the Fermi energy, which is set at $E_{f}=0.42\,\text{eV}$. These edge states originate from three gapped Weyl points (labeled 1–3 in Figure 2(a)) as well as a spin–orbit coupling (SOC)-induced band gap centered at the Fermi level. These gaps belong to bands 55, 56, and 57, with energies of 33.0, 32.0, and 24.0 meV, respectively, with respect to the Fermi energy. We note that a Weyl point is only useful if it is located near the Fermi level, as only states at or near the Fermi energy can exhibit observable effects. Otherwise, contributions from the Weyl nodes may be obscured by other states. In the case of 2D ScV6Sn6, all the three gaps are positioned near the Fermi energy. Notably, Weyl gap 1 is exactly at the Fermi energy, while gaps 2 and 3 are located at the high symmetry K and M points. Our edge state plot in Figure 2(b) shows high intensity in regions originally gapped by SOC in the bulk band structure, indicating electronic states and band crossings. Valence and conduction band crossings at these Weyl points are clearly observed in the edge state band structure (Figure 2(b)). The calculated topological invariant, C, in the absence of external fields, is finite for the three bands, confirming the material as topologically nontrivial with three chiral modes terminating at the edge. Based on this analysis, we classify the ground state of 2D ScV6Sn6 as a Weyl semimetal with edge states near the Fermi energy, due to conduction-valence band crossings at discrete momentum points. pokharel2021electronic ; ortiz2021superconductivity . Since TRS is inherently broken, the stability of the chiral edge states depends solely on the broken TRS and the lattice geometry. The weyl nodes appear in pairs of opposite chirality and are well separated in momentum space to prevent annihilation and ensure their stability. In the presence of SOC, 2D ScV6Sn6 exhibits a quantum anomalous Hall behavior, with the interplay between SOC and magnetic ordering offering additional stability to the chiral modes. For example, Fe3Sn2, a ferromagnetic kagome lattice, exhibits nontrivial topology due to strong SOC and magnetic order yin2018giant . Similarly, 2D ScV6Sn6 retains nontrivial topological characteristics protected by the kagome lattice structure and the strong interplay between SOC and intrinsic magnetic ordering. These features are comparable to those observed in the ferromagnetic Weyl semimetal Co3Sn2S2 bernevig2022progress .